# Chromatic and achromatic numbers of unitary addition Cayley graphs

### v2

@@ 1 Introduction
Throughout this paper, all rings are assumed to be finite rings with unity. There are many different graphs that can be associated to a ring. One may use combinatorial properties of the graph to study algebraic properties of the ring, or perhaps choose the ring so that the graph constructed has some desired property. The recent monograph of Anderson, Asir, Badawi, and Chelvam [3] discusses several graphs obtained from rings such as zero-divisor graphs, total graphs, and Cayley graphs. The subject of this work is unitary addition Cayley graphs and their clique, chromatic and achromatic numbers. These graphs have been the subject of many papers. Their definition is as follows.
@@
Other graph parameters, including girth, Hamiltonicity, and planarity, were studied in [14]. Let us now focus on what is known regarding the clique number, chromatic number, and achromatic number of $\mathcal{U}(\mathbb{Z}_{n})$. Observe that since $\mathcal{U}(\mathbb{Z}_{n})$ is bipartite when $n$ is even, $\chi(\mathcal{U}(\mathbb{Z}_{n}))=2$. Sinha, Garg, and Singh [14] proved that for $p$ an odd prime, and an odd integer $n$ having $m\geq 2$ distinct prime factors,
@@
These same results were obtained by Palanivel and Chithra [10] who also investigated the graphs $\mathcal{U}(\mathbb{Z}_{n})$. For a graph $G$, let $\alpha(G)$ and $\omega(G)$ denote the independence number and clique number of $G$, respectively. Again, since $\mathcal{U}(\mathbb{Z}_{n})$ is bipartite for even $n$, we have $\omega(\mathcal{U}(\mathbb{Z}_{n}))=2$. Palanivel and Chithra [10] proved that for an odd prime $p$ and integer $k\geq 2$ ,
@@
As the lower and upper bounds match, equality holds throughout. The chromatic number of $\mathcal{U}(\mathbb{Z}_{n})$ was studied further by Promsakon [11] who also obtained some of the results of [10]. Thus, the value of $\chi(\mathcal{U}(\mathbb{Z}_{n}))$ has been determined when $n$ is even, and when $n$ is a power of an odd prime. However, when $n$ has more than one prime factor, exact formulas for the clique and chromatic number are not known. Our first theorem gives such formulas for all odd $n$. Furthermore, it applies to a large class of rings that includes $\mathbb{Z}_{n}$ as a special case. Recall that a commutative ring with unity is a local ring if it has a unique maximal ideal. One of the classical results in ring theory is that every finite commutative ring is a product of local rings.
@@
Let $R$ be a finite commutative ring with an odd number of elements that is the direct product of $m$ local rings, say $R=R_{1}\times\cdots\times R_{m}$ where $M_{i}$ is the unique maximal ideal in $R_{i}$. Then
@@
In the case of $\mathbb{Z}_{n}$ with $n$ an odd integer having prime factorization $n=p_{1}^{k^{1}}\cdots p_{m}^{k_{m}}$, we have the isomorphism $\mathbb{Z}_{n}\cong\mathbb{Z}_{p_{1}^{k_{1}}}\times\cdots\times\mathbb{Z}_{p_{m}^{k_{m}}}$. Let us assume $3\leq p_{1}<p_{2}<\dots<p_{m}$. Each factor $\mathbb{Z}_{{p_{i}}^{k_{i}}}$ is a local ring where the unique maximal ideal is $\langle p_{i}\rangle$. Therefore, $|\mathbb{Z}_{{p_{i}}^{k_{i}}}|-|\langle p_{i}\rangle|=p_{i}^{k_{i}}-p_{i}^{k_{i}-1}=\phi(p_{i}^{k_{i}})$. Using the multiplicative property of $\phi$,
@@
Theorem 1.1 also gives the clique and chromatic number of the unitary addition graph of the Gaussian integers modulo $n$, where $n$ is odd, studied by Bahrami and Jahani-Nezhad ([4], Section 3.3).
Originally motivated by a coloring problem on hypercubes, discussed further in Section 5, we also investigated the achromatic number of unitary addition Cayley graphs. The achromatic number of a graph $G$, denoted $\chi_{a}(G)$, is the maximum number of colors in a proper coloring where there is at least one edge between each pair of color classes. In 1967, Harary, Hedetniemi, and Prins introduced these types of colorings in connection with the chromatic number of a graph [8]. Achromatic colorings have been studied extensively since then [5, 6, 7] and in general, they are difficult to find. Not surprisingly, less is known about $\chi_{a}(\mathcal{U}(\mathbb{Z}_{n}))$ in comparison to $\chi(\mathcal{U}(\mathbb{Z}_{n}))$. Momrit and Promsakon [9] proved that for any odd prime $p$ and integer $k\geq 1$,
@@
Thus, the achromatic number of $\mathcal{U}(\mathbb{Z}_{n})$ is known only when $n$ is a power of a prime. In some cases the chromatic number and achromatic number of $\mathcal{U}(\mathbb{Z}_{n})$ are equal. This occurs when $n$ consists of only one prime factor:
@@
When $n$ is not a power of a prime, these values can differ substantially. For example, while $\chi(\mathcal{U}(\mathbb{Z}_{2p}))=2$, we have $\chi_{a}(\mathcal{U}(\mathbb{Z}_{2p}))=p$. In Section 3 we show that $\chi_{a}(\mathcal{U}(\mathbb{Z}_{2p}))=p$, and prove the following lower bound that applies when $n$ is the product of two odd primes.
@@
It could be the case that the coloring proving the lower bound in Theorem 1.3 is best possible. Nevertheless, Theorem 1.4 determines the achromatic number for a non-trivial infinite family of Cayley sum graphs.
Conjecture 1.5.
If $q>p\geq 3$ are primes, then
$\chi_{a}(\mathcal{U}(\mathbb{Z}_{pq}))=\frac{pq+1}{2}.$
@@
Let $R$ be a finite commutative ring with unity having an odd number of elements. Assume that $R$ is a product of local rings, say $R=R_{1}\times\cdots\times R_{m}$. Let $M_{i}$ be the unique maximal ideal in $R_{i}$. It is known that the units of $R_{i}$ are the elements of $R_{i}\backslash M_{i}$, and $\prod_{i=1}^{m}(R_{i}\backslash M_{i})$ is the set of units in $R$. Observe then that if $x=(x_{1},\dots,x_{m})$ and $y=(y_{1},\dots,y_{m})$ are distinct elements in $R$, then $x+y$ is a non-unit if and only if $x_{i}+y_{i}\in M_{i}$ for some $1\leq i\leq m$. In terms of the graph $\mathcal{U}(R_{1}\times\cdots\times R_{m})$, a set of vertices $\mathcal{A}$ forms an independent if for each $x\neq y\in\mathcal{A}$, the sum $x+y$ has at least one coordinate in one of the maximal ideals. We will use this simple fact frequently in our argument.
First we will prove the upper bound
@@
The union $\cup_{i=1}^{m}A_{i}$ is the set of non-units in $R$. Let $i\in\{1,2,\dots,m\}$. For any distinct $x,y\in A_{i}$, the $i$-th coordinate of $x+y$ is in $M_{i}$ so that $x+y$ is not a unit. Therefore, each $A_{i}$ is an independent set.
@@
for two distinct choices of exponents $(i_{1},\dots,i_{m})$ and $(j_{1},\dots,j_{m})$. Without loss of generality, assume $i_{1}=0$ and $j_{1}=1$. Then $x_{1}=-x_{1}$ so that $x_{1}+x_{1}=0_{R_{1}}$. Since $x_{1}\notin M_{1}$, we know that $x_{1}\neq 0_{R_{1}}$ and so $x_{1}$ has additive order 2. This contradicts the fact that $R$ has an odd number of elements. We conclude that $|B|=2^{m}$ and all equivalence classes have $2^{m}$ elements. Let $B_{1},\dots,B_{\tau}$ be the equivalence classes under the relation $\sim$ on $R^{*}$. We will complete the proof by showing that each $B_{i}$ is an independent set. Let $x$ and $y$ be distinct vertices in $B_{i}$. Then
@@
$m+\prod_{i=1}^{m}\left(\frac{|R_{i}|-|M_{i}|}{2}\right)\leq\omega(\mathcal{U}(R_{1}\times\cdots\times R_{m})).$ (2)
Fix an $i$ with $1\leq i\leq m$. Let $\pi_{i}:R_{i}\rightarrow R_{i}/M_{i}$ be the quotient map $\pi_{i}(r)=r+M_{i}$. Since $M_{i}$ is maximal, $R_{i}/M_{i}$ is a finite division ring and thus, is a field (or recall $R$ is commutative). Since $|R|$ is odd, we have the $|R_{i}/M_{i}|=q_{i}$ for some odd prime power $q_{i}$. Let $\hat{S}_{i}$ be a set of $\frac{q_{i}-1}{2}$ non-zero elements of $R_{i}/M_{i}$ such that
@@
for any odd prime $p$, where the lower bound $\chi_{a}(\mathcal{U}(\mathbb{Z}_{2p}))\geq p$ is proved in [9]. ∎
Now we turn our attention to the main part of this section, which is the proof of Theorem 1.4. Let us start with a definition .
@@ 5 Concluding Remarks
An interesting question is given a graph $G$, how close are $\chi_{a}(G)$ and/or $\psi(G)$ to $\sqrt{2e(G)+1/4}+1/2$.
Let $\{G_{n}\}_{n=1}^{\infty}$ be a sequence of graphs such that $|V(G_{n})|\rightarrow\infty$ as $n\rightarrow\infty$. Ahlswede et al. [1] asked for an explicit sequence of graphs $G_{n}$ such that
@@
It is known that the sequence of paths $\{P_{1},P_{2},\dots\}$ and the sequence of cycles $\{C_{3},C_{4},\dots\}$ are almost optimally complete. More generally, any sequence of graphs of bounded degree is almost optimally complete [5]. By Theorem 1.4, $\chi_{a}(\mathcal{U}(\mathbb{Z}_{3q}))=\frac{3q+1}{2}$ for all primes $q\geq 5$. There is roughly a factor of $\frac{8}{3}$ between $\binom{\chi_{a}(\mathcal{U}(\mathbb{Z}_{3q}))}{2}=\frac{9q^{2}-1}{8}$ and $e(\mathcal{U}(\mathbb{Z}_{3q}))=\frac{1}{2}(3q-1)\cdot 2(q-1)=(3q-1)(q-1)$. Therefore, this sequence of graphs is not almost optimally complete.
In light of these questions of Ahlswede et al. [1] and Roichman [12], an interesting sequence to investigate is $\{\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\cdots p_{m}})\}_{m=1}^{\infty}$ where $3=p_{1}<p_{2}<\dots<p_{m}<\dots$ is the sequence of odd primes. We can lower bound the number of edges in $\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}})$ using results in number theory from the 1960’s. Writing $p_{N}$ for the $N$-th prime, the work of Rosser and Shoenfeld [13] gives $p_{N}\leq 2N\log N$ for all $N\geq 6$, and $\prod_{p\leq x}\left(1-\frac{1}{p}\right)>\frac{\textup{exp}(-\gamma)}{\log x}\left(1-\frac{1}{2\log^{2}x}\right)$ for $x\geq 285$. Here the product is over all prime numbers that are less than or equal to $x$, $\gamma$ is the Euler-Mascheroni constant, and $\textup{exp}(t):=e^{t}$. Using these inequalities we have for large enough $m$,
@@
This leads us to ask whether or not $\psi(\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}}))$ is $o(\frac{p_{1}p_{2}\cdots p_{m}}{\sqrt{\log(m\log m)}})$? Since a pseudo-achromatic coloring need not be proper, proving a good upper bound on $\psi(\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}}))$ could be more difficult than proving a good upper bound on $\chi_{a}(\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}}))$. In this paper all colorings considered were proper, and this assumption was used in the proof of Theorem 1.4. We close by noting that a resolution of Conjecture 1.5 and studying more prime factors could lead to ideas for answering the following interesting question.
Question 5.1.
If $3\leq p_{1}<p_{2}<\dots<p_{m}$ is the sequence of odd primes, then is it true that
$\chi_{a}(\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\cdots p_{m}}))=O\left(\frac{p_{1}p_{2}\cdots p_{m}}{m}\right)~{}?$

### v3

@@ 1 Introduction
Throughout this paper, all rings are assumed to be finite rings with unity. There are many different graphs that can be associated to a ring. One may use combinatorial properties of the graph to study algebraic properties of the ring, or perhaps choose the ring so that the graph constructed has some desired property. The recent monograph of Anderson, Asir, Badawi, and Chelvam [3] discusses several graphs obtained from rings such as zero-divisor graphs, total graphs, and Cayley graphs. The subject of this work is unitary addition Cayley graphs and their clique, chromatic and achromatic numbers. These graphs have been the subject of many papers [2, 4, 5, 10, 11, 12, 13, 14, 17]. Their definition is as follows.
@@
In 2007, Klotz and Sander [10] proved a number of results about the graphs $\mathcal{U}(\mathbb{Z}_{n})$. Since their work, many graphs parameters such as the girth, Hamiltonicity, planarity, and the spectrum of $\mathcal{U}(\mathbb{Z}_{n})$ have been studied. Let us now focus on what is known regarding the clique number, chromatic number, and achromatic number of $\mathcal{U}(\mathbb{Z}_{n})$. Observe that $\chi(\mathcal{U}(\mathbb{Z}_{n}))=2$ since $\mathcal{U}(\mathbb{Z}_{n})$ is bipartite when $n$ is even. Sinha, Garg, and Singh [17] proved that for $p$ an odd prime and an odd integer $n$ having $m\geq 2$ distinct prime factors,
@@
These same results were obtained by Palanivel and Chithra [13] who also investigated $\mathcal{U}(\mathbb{Z}_{n})$. For a graph $G$, let $\alpha(G)$ and $\omega(G)$ denote the independence number and clique number of $G$, respectively. Again, since $\mathcal{U}(\mathbb{Z}_{n})$ is bipartite when $n$ is even, we have $\omega(\mathcal{U}(\mathbb{Z}_{n}))=2$. Palanivel and Chithra [13] proved that for an odd prime $p$ and integer $k\geq 2$,
@@
As the lower and upper bounds match, equality holds throughout. The chromatic number of $\mathcal{U}(\mathbb{Z}_{n})$ was studied further by Promsakon [14] who also obtained some of the results of [13]. Thus, the value of $\chi(\mathcal{U}(\mathbb{Z}_{n}))$ has been determined when $n$ is even, and when $n$ is a power of an odd prime. However, when $n$ has more than one prime factor exact formulas for the clique and chromatic number are not known. Our first theorem gives such formulas for all odd $n$. Furthermore, it applies to a large class of rings that includes $\mathbb{Z}_{n}$ as a special case. Recall that a commutative ring with unity is a local ring if it has a unique maximal ideal. One of the classical results in ring theory is that every finite commutative ring is a product of local rings.
@@
Let $R$ be a finite commutative ring with an odd number of elements that is the direct product of $m$ local rings, say $R=R_{1}\times\cdots\times R_{m}$ where $M_{i}$ denotes the unique maximal ideal in $R_{i}$. Then
@@
In the case of $\mathbb{Z}_{n}$ with $n$ an odd integer having prime factorization $n=p_{1}^{k^{1}}\cdots p_{m}^{k_{m}}$, we have the ring isomorphism $\mathbb{Z}_{n}\cong\mathbb{Z}_{p_{1}^{k_{1}}}\times\cdots\times\mathbb{Z}_{p_{m}^{k_{m}}}$. Let us assume that $3\leq p_{1}<p_{2}<\dots<p_{m}$. Each factor $\mathbb{Z}_{{p_{i}}^{k_{i}}}$ is a local ring where the unique maximal ideal is $\langle p_{i}\rangle$. Therefore, $|\mathbb{Z}_{{p_{i}}^{k_{i}}}|-|\langle p_{i}\rangle|=p_{i}^{k_{i}}-p_{i}^{k_{i}-1}=\phi(p_{i}^{k_{i}})$. Using the multiplicative property of $\phi$,
@@
The previously best known bounds on the clique and chromatic number on $\mathcal{U}(\mathbb{Z}_{n})$, as well as our contributions, are summarized in the following table. In rows 2, 3, 4, and 5, $p,p_{1},\dots,p_{m}$ are odd primes with $3\leq p_{1}<p_{2}<\dots<p_{m}$ and $k$ is a positive integer.
number of vertices $\omega$ $\chi$ Reference
$n$ where $n$ is even $2=\omega(\mathcal{U}(\mathbb{Z}_{n}))$ $\chi(\mathcal{U}(\mathbb{Z}_{n}))=2$ $\mathcal{U}(\mathbb{Z}_{n})$ is bipartite
$p$ $\frac{p+1}{2}=\omega(\mathcal{U}(\mathbb{Z}_{p}))$ $\chi(\mathcal{U}(\mathbb{Z}_{p}))=\frac{p+1}{2}$ [14, 17]
$p^{k}$ $\frac{\phi(p^{k})}{2}+1=\omega(\mathcal{U}(\mathbb{Z}_{p^{k}}))$ $\chi(\mathcal{U}(\mathbb{Z}_{p^{k}}))=\frac{\phi(p^{k})}{2}+1$ [13, 14]
$n=p_{1}^{k_{1}}\cdots p_{m}^{k_{m}}$ $\frac{p_{1}+1}{2}\leq\omega(\mathcal{U}(\mathbb{Z}_{n}))$ $\chi(\mathcal{U}(\mathbb{Z}_{n}))\leq\frac{\phi(n)}{2^{m}}+m$ [13, 14, 17]
$n=p_{1}^{k_{1}}\cdots p_{m}^{k_{m}}$ $\frac{\phi(n)}{2^{m}}+m=\omega(\mathcal{U}(\mathbb{Z}_{n}))$ $\chi(\mathcal{U}(\mathbb{Z}_{n}))=\frac{\phi(n)}{2^{m}}+m$ Corollary 1.2
Current Bounds on $\omega(\mathcal{U}(\mathbb{Z}_{n})$ and $\chi(\mathcal{U}(\mathbb{Z}_{n})$ and Corollary 1.2
Other unitary graphs, aside from $\mathcal{U}(\mathbb{Z}_{n})$ have been investigated. For example, Theorem 1.1 gives the clique and chromatic number of the unitary addition graphs of the Gaussian integers modulo $n$, where $n$ is odd. These graphs were studied by Bahrami and Jahani-Nezhad ([4], Section 3.3).
Originally motivated by a coloring problem on hypercubes, discussed in further detail in Section 5, we also investigated the achromatic number of unitary addition Cayley graphs. The achromatic number of a graph $G$, denoted $\chi_{a}(G)$, is the maximum number of colors in a proper coloring where there is at least one edge between each pair of color classes. In 1967, Harary, Hedetniemi, and Prins introduced these types of colorings in connection with the chromatic number of a graph [9]. Achromatic colorings have been studied extensively since then [6, 7, 8]. In general, they are difficult to find. Not surprisingly, less is known about $\chi_{a}(\mathcal{U}(\mathbb{Z}_{n}))$ compared to $\chi(\mathcal{U}(\mathbb{Z}_{n}))$. Momrit and Promsakon [12] proved that for any odd prime $p$ and integer $k\geq 1$,
@@
Thus, the achromatic number of $\mathcal{U}(\mathbb{Z}_{n})$ is known only when $n$ is a power of a prime. In some cases, the chromatic number and the achromatic number of $\mathcal{U}(\mathbb{Z}_{n})$ are equal. This occurs when $n$ consists of only one prime factor:
@@
When $n$ is not a prime power, these values can differ substantially. For example, while $\chi(\mathcal{U}(\mathbb{Z}_{2p}))=2$, we have $\chi_{a}(\mathcal{U}(\mathbb{Z}_{2p}))=p$. In Section 3 we show that $\chi_{a}(\mathcal{U}(\mathbb{Z}_{2p}))=p$ and prove the following lower bound that applies when $n$ is the product of two odd primes.
@@
Theorem 1.4 determines the achromatic number for a non-trivial infinite family of Cayley sum graphs. Initially it was thought that the formula in Theorem 1.4 could hold more generally and we conjectured that the achromatic number of $\mathcal{U}(\mathbb{Z}_{pq})$ is $\frac{pq+1}{2}$ for $q>p\geq 3$ prime. If this were true, then it would show that the bound from Theorem 1.3 is best possible. In addressing questions raised by the referee report(s), we discovered that this conjecture is false. Using a computer search, we were able to find an achromatic coloring of $\mathcal{U}(\mathbb{Z}_{35})$ with 19 colors, exceeding $\frac{5\cdot 7+1}{2}$. An achromatic coloring of $\mathcal{U}(\mathbb{Z}_{55})$ with $30>\frac{5\cdot 11+1}{2}$ colors was also found (the authors would like to thank the reviewer(s) for comments regarding our conjecture).
@@
Let $R$ be a finite commutative ring with unity having an odd number of elements. Assume that $R$ is a product of local rings, say $R=R_{1}\times\cdots\times R_{m}$. Let $M_{i}$ be the unique maximal ideal in $R_{i}$. It is known that the units of $R_{i}$ are the elements of $R_{i}\backslash M_{i}$, and $\prod_{i=1}^{m}(R_{i}\backslash M_{i})$ is the set of units in $R$. Observe then that if $x=(x_{1},\dots,x_{m})$ and $y=(y_{1},\dots,y_{m})$ are distinct elements in $R$, then $x+y$ is a non-unit if and only if $x_{i}+y_{i}\in M_{i}$ for some $1\leq i\leq m$. In terms of the graph $\mathcal{U}(R_{1}\times\cdots\times R_{m})$, a set of vertices $\mathcal{A}$ forms an independent set if for each $x\neq y\in\mathcal{A}$, the sum $x+y$ has at least one coordinate in one of the maximal ideals. We will use this fact frequently in our argument.
First we prove the upper bound
@@
Note that every $x\in A_{i}$ has its $i$-th coordinate in $M_{i}$, thus such an $x$ is a non-unit in $R$. Further, every non-unit in $R$ has a coordinate in $M_{j}$ for some minimal $j$ and must be an element of $A_{j}$. It follows that the union $\cup_{i=1}^{m}A_{i}$ is the set of non-units in $R$. Let $i\in\{1,2,\dots,m\}$. For any distinct $x,y\in A_{i}$, the $i$-th coordinate of $x+y$ is in $M_{i}$ so that $x+y$ is not a unit. Therefore, each $A_{i}$ is an independent set.
@@
for two distinct choices of exponents $(i_{1},\dots,i_{m})$ and $(j_{1},\dots,j_{m})$. Without loss of generality, assume $i_{1}=0$ and $j_{1}=1$. Then $x_{1}=-x_{1}$ so that $x_{1}+x_{1}=0_{R_{1}}$. Since $x_{1}\notin M_{1}$, we know that $x_{1}\neq 0_{R_{1}}$ and so $x_{1}$ has additive order 2 in $R_{1}$. This contradicts the fact that $R$ has an odd number of elements. We conclude that $|B|=2^{m}$ and all equivalence classes have $2^{m}$ elements. Let $B_{1},\dots,B_{\tau}$ be the equivalence classes under the relation $\sim$ on $R^{*}$. We will complete the proof by showing that each $B_{i}$ is an independent set. Let $x$ and $y$ be distinct vertices in $B_{i}$. Then
@@
$m+\prod_{i=1}^{m}\left(\frac{|R_{i}|-|M_{i}|}{2}\right)\leq\omega(\mathcal{U}(R_{1}\times\cdots\times R_{m}))$ (2)
by demonstrating a clique of size $m+|R^{*}|/2^{m}$. Fix an $i$ with $1\leq i\leq m$. Let $\pi_{i}:R_{i}\rightarrow R_{i}/M_{i}$ be the quotient map $\pi_{i}(r)=r+M_{i}$. Since $M_{i}$ is maximal, $R_{i}/M_{i}$ is a finite division ring and thus, is a field (or recall $R$ is commutative). Since $|R|$ is odd, we have the $|R_{i}/M_{i}|=q_{i}$ for some odd prime power $q_{i}$. Let $\hat{S}_{i}$ be a set of $\frac{q_{i}-1}{2}$ non-zero elements of $R_{i}/M_{i}$ such that
@@
for any odd prime $p$ (the lower bound $\chi_{a}(\mathcal{U}(\mathbb{Z}_{2p}))\geq p$ is proved in [12]). ∎
Now we turn our attention to the main part of this section which is the proof of Theorem 1.4. Let us start with a definition.
@@ 5 Concluding Remarks
An interesting question is given a graph $G$, how close are $\chi_{a}(G)$ and $\psi(G)$ to $\sqrt{2e(G)+1/4}+1/2$. Let $\{G_{n}\}_{n=1}^{\infty}$ be a sequence of graphs such that $|V(G_{n})|\rightarrow\infty$ as $n\rightarrow\infty$. Ahlswede et al. [1] asked for an explicit sequence of graphs $G_{n}$ such that
@@
It is known that the sequence of paths $\{P_{1},P_{2},\dots\}$ and the sequence of cycles $\{C_{3},C_{4},\dots\}$ are almost optimally complete. More generally, any sequence of graphs of bounded degree is almost optimally complete [6]. By Theorem 1.4, $\chi_{a}(\mathcal{U}(\mathbb{Z}_{3q}))=\frac{3q+1}{2}$ for all primes $q\geq 5$. There is roughly a factor of $\frac{8}{3}$ between $\binom{\chi_{a}(\mathcal{U}(\mathbb{Z}_{3q}))}{2}=\frac{9q^{2}-1}{8}$ and $e(\mathcal{U}(\mathbb{Z}_{3q}))=\frac{1}{2}(3q-1)\cdot 2(q-1)=(3q-1)(q-1)$. Therefore, the sequence $\{\mathcal{U}(\mathbb{Z}_{3q}):q=5,7,11,\dots\}$ is not almost optimally complete.
In light of these questions of Ahlswede et al. [1] and Roichman [15], an interesting sequence to investigate is $\{\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\cdots p_{m}})\}_{m=1}^{\infty}$ where $3=p_{1}<p_{2}<\dots<p_{m}<\dots$ is the sequence of odd primes. We can lower bound the number of edges in $\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}})$ using results in number theory. Writing $p_{N}$ for the $N$-th prime, the work of Rosser and Shoenfeld [16] gives $p_{N}\leq 2N\log N$ for all $N\geq 6$, and $\prod_{p\leq x}\left(1-\frac{1}{p}\right)>\frac{\textup{exp}(-\gamma)}{\log x}\left(1-\frac{1}{2\log^{2}x}\right)$ for $x\geq 285$. Here the product is over all prime numbers that are less than or equal to $x$, $\gamma$ is the Euler-Mascheroni constant, and $\textup{exp}(t):=e^{t}$. Using these inequalities we have for large enough $m$,
@@
This leads us to ask whether or not $\psi(\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}}))$ is $o(\frac{p_{1}p_{2}\cdots p_{m}}{\sqrt{\log(m\log m)}})$? Since a pseudo-achromatic coloring need not be proper, proving a good upper bound on $\psi(\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}}))$ could be more difficult than proving a good upper bound on $\chi_{a}(\mathcal{U}(\mathbb{Z}_{p_{1}p_{2}\dots p_{m}}))$. In this paper all colorings considered were proper and this assumption was used in the proof of Theorem 1.4.